OpenStudy (anonymous):

solve for x and y 12x + y = −35 12x + 6y = −3

OpenStudy (bahrom7893):

12x = -35-y -35-y+6y=-3

OpenStudy (bahrom7893):

-35 + 5y = =-3 5y = 32 y = 32/5

OpenStudy (bahrom7893):

12x = -35 - 32/5 12x = (-175-32)/5 = -207/5

OpenStudy (bahrom7893):

uhmm can someone confirm? the fractions look ugly... maybe i made a mistake